OpenStudy (karinamtz0213):

Please help Verify: Sec^2x(1-cos^2x)=tan^2x

OpenStudy (anonymous):

sec^2x (1-cos^2x) = tan^2x sec^2x = 1/cos^2x 1-cos^2x = sin^2x Therefore: 1/cos^2x (sin^2x) = sin^2x/cos^2x = tan^2x

OpenStudy (karinamtz0213):

Thanks alot.